OpenStudy (anonymous):

Separate 72 into two parts, one of which shall be 1/3 of the other.

OpenStudy (anonymous):

1/3a + a=72

OpenStudy (anonymous):

How about this? x = smaller number 3x = larger number 4x = 72 4x/4 = 72/4 x = 18

OpenStudy (anonymous):

that works